# Edge AI: A Taxonomy, Systematic Review and Future Directions

### v1

@@ VIII-I Heterogeneity
A multi-prototype-based federated learning enhances model inference by leveraging multiple weighted prototypes rather than relying on a single prototype, which can be incomplete and ambiguous [90]. This can aid in calculating these prototypes locally using clustering algorithms like k-means, ensuring effective representation of different client data distributions. Aggregating these multiple prototypes improves robustness against non-IID data distributions, leading to higher test accuracy and more efficient communication. This approach demonstrates significant improvements in accuracy and convergence rates, making it a promising direction for handling heterogeneity in Edge AI.
@@ VIII-J Security
Utilizing AI-based quantum-safe cybersecurity automation, improving device and sensor security, and integrating quantum-safe solutions to counteract the threats posed by advanced quantum computing [158]. The research will focus on developing scalable and efficient cybersecurity systems, including AI-driven automation for threat detection and mitigation and using blockchain for secure communications [159]. Additionally, creating robust test environments for cybersecurity validation will ensure the effectiveness of these solutions in diverse operational scenarios [160]. These developments aim to provide a comprehensive security framework for future edge AI systems, ensuring resilience against evolving cyber threats.
@@ IX Summary and Conclusions
The systematic review analysis of Edge AI provides a comprehensive overview of the present status of research in edge intelligence and its applications. The significance of these findings lies in the need for Edge AI systems to consider infrastructure, resource management, and the scale of ML models. According to the findings of the study, it is essential to conduct a thorough examination of both the positive and negative aspects of prior research in order to identify any potential gaps in research and to estimate prospective developments and concerns.
The study emphasises the importance of using a systematic approach to record and evaluate the existing research in the field of Edge AI. Moreover, it highlights the importance of implementing a standardised procedure to reduce the possible impact of discrepancies in the study. The results of this review have the capacity to ignite a new field of investigation in Edge AI and offer direction for prospective research in this field. The exhaustive examination of Edge AI offers profound insight into the most current study on edge intelligence and its realistic implementations. Moreover, this emphasises the importance of gaining additional understanding about the key factors that govern the choice of Edge AI infrastructure, as well as the effect of the scale of the model used for ML on efficiency and resource allocation.
To summarize, the comprehensive study on Edge AI is to fully evaluate the various AI methodologies. This study integrates all the feasible methodologies incorporates in edge intelligence or AI at the edge. This review is to examine the crucial factors that impact the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, and assess their impact on application efficacy and resource utilization. Furthermore, it investigates the influence of the size of a ML model on the efficacy and resource usage of an Edge AI application. A total of 78 studies have been chosen for this evaluation due to their specific focus on the application of AI in edge computing. In order to enhance our comprehension in the context of AI applied to edge computing, these research were categorized into multiple domains, including infrastructure, resource management, and ML model sizing, among others. Resource supply, allocation, scheduling, and job deployment are crucial considerations in the field of Edge AI.
Acknowledgements

### v2

@@ VIII-I Heterogeneity
In edge AI environments, heterogeneity refers to the diverse nature of data, devices, and communication networks across edge nodes. This diversity introduces challenges in federated learning, where edge devices typically possess non-identically distributed (non-IID) data, varied computational capacities, and inconsistent communication bandwidths. As traditional FL approaches often rely on a single global model, they may struggle to capture the diverse patterns and distributions present across different edge devices. To address these challenges, multi-prototype-based federated learning [208] has emerged as a promising approach for enhancing model inference by leveraging multiple weighted prototypes rather than relying on a single prototype, which can be incomplete and ambiguous [94].
A key aspect of this approach involves calculating local prototypes at each edge device, ensuring that the diverse distributions of client data are effectively represented. These prototypes capture the unique characteristics of each client’s data distribution, enabling a more nuanced aggregation process during global model updates. Clustering algorithms like k-means [208] can be employed locally at each edge device to generate multiple prototypes that correspond to different data clusters within the client’s dataset. By calculating prototypes that reflect the underlying structure of local data, this approach enhances the model’s ability to generalize across heterogeneous client distributions. Furthermore, these multiple weighted prototypes provide a richer representation of client data compared to traditional single-prototype methods, which often oversimplify local data distributions.
Once local prototypes are calculated, these prototypes are aggregated across devices during the global model update process. Rather than averaging model updates from each device, as in traditional FL approaches, multi-prototype-based FL aggregates the prototypes in a weighted manner, where each prototype’s contribution is proportional to the importance of the corresponding data cluster. This approach improves robustness against non-IID [209] data distributions by ensuring that the global model is not overly influenced by outliers or overrepresented data points. Instead, the weighted aggregation process captures the full diversity of data across the edge devices, leading to higher test accuracy and better generalization across all devices.
An important challenge in this approach is ensuring communication efficiency, especially in bandwidth-constrained edge environments. By reducing the need to transmit the entire local model or large datasets, multi-prototype-based FL [210] minimizes communication overhead. Instead of sharing raw data or full model updates, devices only transmit the prototypes and their associated weights. This reduces the amount of information exchanged between the server and the devices, while still enabling effective global model updates. Quantization techniques can also be applied to further reduce the size of transmitted prototypes, allowing for more efficient communication without compromising model performance.
In addition to improving communication efficiency, multi-prototype-based FL demonstrates significant improvements in both accuracy and convergence rates. The ability to capture more representative patterns from local data allows the global model [210] to converge faster, particularly in heterogeneous environments where traditional FL methods tend to suffer from slow convergence due to non-IID data. The richer and more representative model built through multi-prototype aggregation helps the system achieve higher accuracy in a shorter period, reducing the number of communication rounds required to achieve an optimal model.
This approach demonstrates significant improvements in accuracy and convergence rates, making it a promising direction for handling heterogeneity in edge AI. The multi-prototype strategy opens new avenues for edge AI applications where data heterogeneity is a major bottleneck, such as in personalized healthcare, where patient data varies significantly across different locations, or in smart city infrastructures, where sensor data may differ drastically based on environmental conditions. As future research progresses, integrating other advanced clustering techniques like Gaussian Mixture Models (GMMs) [211] or spectral clustering could further enhance the capability of multi-prototype FL systems, enabling even better handling of non-IID data and improving performance in highly heterogeneous environments.
@@ VIII-J Security
As edge AI becomes more prevalent, the security of edge devices, data, and communications becomes a critical concern, particularly with the growing threat of quantum computing, which can break classical encryption methods. One of the primary directions for securing edge AI systems is the integration of AI-based quantum-safe cybersecurity automation. Quantum-safe cryptographic techniques are essential for protecting sensitive data from the computational power of quantum computers, which could easily break traditional public-key encryption systems. Algorithms such as lattice-based cryptography, hash-based signatures, and code-based cryptography are being researched as quantum-resistant alternatives that can secure edge AI devices and communications against quantum attacks [212]. These quantum-safe solutions ensure that even as quantum computing capabilities advance, edge AI systems remain resilient against cryptographic threats.
Improving device and sensor security is another critical focus area. Edge devices, by their nature, are distributed and often deployed in insecure environments, making them vulnerable to physical tampering and cyberattacks. Integrating AI-based security mechanisms that can detect abnormal behavior at the device level is essential for enhancing the security of edge networks. For example, deep learning-based intrusion detection systems (IDS) can be implemented at the edge to monitor incoming traffic for potential threats, such as denial-of-service (DoS) attacks or unauthorized access attempts. These IDS systems can utilize anomaly detection algorithms, such as autoencoders or generative adversarial networks (GANs), to identify deviations from normal traffic patterns and flag potential security breaches in real time. Such AI-driven systems can adapt over time, learning from new attack vectors and updating their detection models to address emerging threats. Additionally, lightweight blockchain-based solutions can be integrated to ensure the secure exchange of data between edge devices by creating immutable records of transactions, further reducing the risk of tampering [213].
In the context of quantum-safe solutions, edge AI systems must also adopt post-quantum cryptographic algorithms to ensure long-term data security. Post-quantum cryptography (PQC) focuses on developing encryption algorithms that are resistant to both classical and quantum attacks. Integrating PQC into edge AI devices ensures that secure communications are maintained even when quantum computers become widely available. Furthermore, using AI to optimize the implementation of PQC algorithms [214], such as by reducing their computational overhead, can make these solutions more practical for deployment in resource-constrained edge environments. Research is also focusing on quantum key distribution (QKD) [215], a technique that leverages quantum mechanics to generate provably secure cryptographic keys. QKD can be used to secure communication between edge devices and central servers, ensuring that keys cannot be intercepted or tampered with, even by quantum adversaries.
The research will focus on developing scalable and efficient cybersecurity systems, including AI-driven automation for threat detection and mitigation and using blockchain for secure communications [213]. A key challenge in securing edge AI is the need for scalability. As the number of connected devices in edge networks increases, cybersecurity solutions must be able to scale to protect millions of devices without introducing significant latency or computational overhead. AI-driven security systems can help address this challenge by automating threat detection and response processes. For example, machine learning models can be trained to detect anomalies in device behavior or network traffic, identifying potential cyberattacks before they can cause damage. In addition to intrusion detection, AI can also automate patch management by identifying vulnerabilities in edge devices and applying security updates in real time, ensuring that devices remain protected against known threats.
Additionally, creating robust test environments for cybersecurity validation will ensure the effectiveness of these solutions in diverse operational scenarios [216]. Test environments, such as digital twins of edge networks, can be used to simulate cyberattacks and validate the effectiveness of AI-driven security solutions. By replicating real-world conditions, these environments enable researchers to fine-tune their algorithms and improve the resilience of edge AI systems. For example, by simulating distributed denial-of-service (DDoS) attacks on a digital twin of an edge network [217], AI-based security systems can be stress-tested and adjusted to ensure their ability to respond to large-scale cyber threats in real time. Furthermore, AI-based systems can be used to predict potential vulnerabilities in edge networks by analyzing historical data and identifying patterns that could lead to future security breaches. This proactive approach to security helps to mitigate risks before they can be exploited by attackers.
These developments aim to provide a comprehensive security framework for future edge AI systems, ensuring resilience against evolving cyber threats. The future of edge AI security lies in the combination of quantum-safe cryptography [218], AI-driven threat detection, and automation. These technologies will allow edge AI systems to remain secure in the face of both classical and quantum-based cyber threats, ensuring that they can operate reliably in increasingly complex and hostile environments. As the number of connected devices in smart cities, autonomous vehicles, and industrial IoT grows, maintaining robust security at the edge will be critical to safeguarding sensitive data and ensuring the integrity of AI-driven processes.
@@ IX Summary and Conclusions
The systematic review analysis of Edge AI provides a comprehensive overview of the present status of research in edge intelligence and its applications. The significance of these findings lies in the need for Edge AI systems to consider infrastructure, resource management, and the scale of ML models. According to the findings of the study, it is essential to conduct a thorough examination of both the positive and negative aspects of prior research to identify any potential research gaps and to estimate prospective developments and concerns.
The study emphasises the importance of using a systematic approach to record and evaluates the existing research in the field of Edge AI. Moreover, it highlights the importance of implementing a standardized procedure to reduce the possible impact of discrepancies in the study. The results of this review have the capacity to ignite a new field of investigation in Edge AI and offer direction for prospective research in this field. The exhaustive examination of Edge AI offers profound insight into the most current study on edge intelligence and its realistic implementations. Moreover, this emphasises the importance of gaining additional understanding about the key factors that govern the choice of Edge AI infrastructure, as well as the effect of the scale of the model used for ML on efficiency and resource allocation.
To summarize, the comprehensive study on Edge AI is to fully evaluate the various AI methodologies. This study integrates all the feasible methodologies incorporated in edge intelligence or AI at the edge. This review examines the crucial factors that impact the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, and assesses their impact on application efficacy and resource utilization. Furthermore, it investigates the influence of the size of an ML model on the efficacy and resource usage of an Edge AI application. A total of 78 studies have been chosen for this evaluation due to their specific focus on the application of AI in edge computing. In order to enhance our comprehension in the context of AI applied to edge computing, these studies were categorized into multiple domains, including infrastructure, resource management, and ML model sizing, among others. Resource supply, allocation, scheduling, and job deployment are crucial considerations in the field of Edge AI.
IX-A Open Challenges
Following facts can be further concluded to improve this survey:
  • EdgeAI Infrastructure Optimization: Future studies can examine Edge, Fog and Cloud systems that can be integrated into EdgeAI to create a hybrid model. In this way, scalable infrastructure solutions for EdgeAI systems can be discussed in detail. The research can focus on optimization techniques for resource allocation and latency in systems with varying workloads.
  • Security and Privacy in EdgeAI: Considering the heterogeneous structure of the nodes that make up EdgeAI systems, security and privacy issues arise for sensitive data (biometric data). Future research can examine the measures and technical methods taken to ensure the security and privacy of data.
  • EdgeAI Applications: Future studies can examine real-world EdgeAI applications such as smart cities and IoT-based healthcare systems. In this way, application challenges and solutions can be provided for researchers.
Declarations
  • Funding
@@ IX-A Open Challenges
  • Conflict of interest/Competing interests
    On behalf of all authors, the corresponding author states that there is no conflict of interest.
  • Ethics approval
    Not Available
  • Consent to participate
    Not Available
  • Consent for publication
    Not Available
  • Availability of data and materials
    Not Available
  • Code availability
    Not Available
  • Authors’ contributions
    Sukhpal Singh Gill (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Muhammed Golec (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Jianmin Hu (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Minxian Xu (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Junhui Du (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Huaming Wu (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Guneet Kaur Walia (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Software: Lead; Validation: Lead; Writing – original draft: Lead) Subramaniam Subramanian Murugesan (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Babar Ali (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Writing – original draft: Lead) Mohit Kumar (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Kejiang Ye (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Prabal Verma (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Surendra Kumar (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Felix Cuadrado (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead) Steve Uhlig (Conceptualization: Lead; Data curation: Lead; Formal analysis: Lead; Investigation: Lead; Methodology: Lead; Validation: Lead; Writing – original draft: Lead)